# Dahyeon Kye, Changhyun Roh, Sukhun Ko, Chanho Eom, and Jihyong Oh https://github.com/CMLab-Korea/Awesome-Video-Frame-Interpolation

### v1

@@ 4.2 Occlusion
Achieving high-quality (HQ) interpolation demands accurate motion estimation as well as a proper understanding of occlusions. Otherwise, severe artifacts are likely to appear in the interpolated frames, particularly near motion boundaries. For two consecutive input frames, certain pixels in the intermediate frame may not correspond to any observable region in either input, creating ambiguity in determining the correct content for these occluded regions [188]. As shown in Fig. 10 (b), such occlusions can occur when previously hidden objects become visible or when objects move toward the camera, revealing regions that were not seen in either input. Naively blending warped inputs often results in severe artifacts, most notably ghosting artifacts [87], where an object is not only incorrectly projected from its previous location but also appears as a duplicate at its correct position due to the lack of sufficient visual cues. This is especially problematic in disoccluded regions, areas newly revealed in the intermediate frame but absent in both inputs, such as when an object emerges from behind another or moves directly toward the viewpoint. In these cases, the absence of visual evidence introduces ambiguity, making it unclear what content should be synthesized. To resolve this, modern VFI methods incorporate explicit occlusion reasoning to guide the synthesis process.
@@
Overall, occlusion-aware VFI remains a critical challenge, particularly in dynamic scenes with depth discontinuities or disoccluded motion. As such, SOTA models increasingly combine multiple strategies, such as masking, depth priors, feature similarity, or forward-backward consistency [184, 87] to recover plausible content in ambiguous regions and maintain temporal coherence in the output.
4.3 Lighting Variation
@@
4.4 Non-linear Motion
Many early VFI methods [16, 5, 30, 33, 7, 73, 76, 74, 6, 35, 60, 85] assume linear or uniform motion between input frames. Under this assumption, objects move along straight trajectories at constant velocity, allowing motion estimation based on simple temporal interpolation. Flow-based [16, 5, 33, 6], kernel-based [33], and even phase-based models [49] often rely on this assumption implicitly. However, in real-world scenarios, motion is frequently non-linear due to acceleration, deceleration, or directional change. As shown in Fig. 10 (d), a sliding ball accelerates along a curved path, violating the linear motion prior and introducing significant estimation error.
To address these limitations, researchers have proposed higher-order motion modeling that extends beyond linear assumptions. Since most existing methods operate on only two input frames, they are inherently under-constrained and forced to assume simple motion. To overcome this, several methods incorporate multiple input frames (typically four) to capture richer temporal variations and better approximate non-linear motion. QVI [76] introduces a quadratic motion model that fits second-order trajectories over four consecutive input frames. Specifically, it takes $(I_{-1},I_{0},I_{1},I_{2})$ as inputs and predicts an intermediate frame $I_{t}$ for arbitrary $t\in(0,1)$. By modeling both velocity and acceleration from surrounding frames, QVI enables the network to better handle curved or time-varying motion paths. This parametric formulation allows the model to explicitly account for motion curvature. EQVI [77] further refines by combining offset-based warping with temporal embeddings, improving precision and robustness under complex motions. More recently, IQ-VFI [86] introduces an implicit motion representation using a coordinate-based MLP that adapts to arbitrary motion patterns without requiring predefined trajectory assumptions. These works collectively emphasizes the importance of modeling non-linear motion directly, especially in multi-frame settings. However, these simple mathematical models cannot completely capture the complexities and irregularities of real-world motions.
As the field progresses, a new challenge velocity ambiguity [90] appears. When only two frames are available, multiple plausible motion trajectories can explain the observed displacement, making the underlying motion inherently under-constrained. This ambiguity becomes especially pronounced in scenes involving curved motion or directional switches, such as bouncing balls or rotating limbs. To tackle this, Zhong et al. [90] introduces a velocity embedding module that learns to disambiguate temporal dynamics by jointly reasoning over motion direction and temporal consistency. It separates appearance modeling from motion estimation, which enhances robustness in complex scenes. BiM-VFI [137] takes a complementary perspective by designing an explicit bidirectional motion descriptor. Its Bidirectional Motion Fields (BiM) encode angular and magnitude differences relative to the intermediate time, enabling accurate modeling of curved, asymmetric, or velocity-changing trajectories. BiM-VFI further integrates these representations into a BiM-guided flow estimator and motion-aware refinement network, yielding temporally coherent results in non-linear regimes. These recent advances signal a shift from rigid linear motion priors toward flexible, context-aware motion modeling. By extending temporal supervision and refining motion representations, either through quadratic formulations, implicit embeddings, or directional velocity fields, modern VFI methods now offer significantly improved performance in complex motion scenarios that were previously underexplored.
5 Datasets and Evaluation
5.1 Datasets
To facilitate training and evaluation across varying temporal resolutions and motion complexities, numerous VFI datasets have been developed. Table I provides a high-level summary of commonly used datasets categorized into triplet and multi-frame types. We describe each dataset in detail below.
5.1.1 Triplet Datasets
Early learning-based VFI approaches primarily rely on triplet datasets, where two input frames are used to predict the temporally centered GT frame. This configuration aligns with CTFI settings (Sec. 3.1). Some datasets are further extended to seven-frame sequences [6] for evaluating frame-rate upsampling.
  • Middlebury [189]: Originally designed for optical flow, Middlebury contains short video clips with moderate complexity. Its small size limits scalability, but it remains a standard benchmark for consistency evaluation.
  • UCF101 [178, 16]: A human action dataset from which a small subset of triplets is used for VFI. Due to its LR and simple motion, it is mainly used for training or sanity checks.
  • Vimeo90K [6]: A widely adopted benchmark with diverse scenes and consistent format. It offers clean supervision and balanced motion complexity, making it ideal for comparative analysis.
  • SNU-FILM [50]: Constructed from high-speed footage and categorized by motion difficulty, SNU-FILM enables evaluation across varying levels of motion, occlusion, and blur.
  • ATD-12K [180]: A large-scale animation dataset with rich stylistic diversity. Its variation in artistic textures and motion patterns supports both general-purpose and domain-specific evaluation.
5.1.2 Multi-frame Datasets
TABLE I: Summary and comparison of popular datasets for VFI.
The dataset types
@@ 5.1.2 Multi-frame Datasets
represents Triplet dataset,
@@
represents Multi-frame dataset.
@@
Multi-frame datasets enable dense temporal supervision and are commonly used in both CTFI and ATFI (Sec. 3.2) settings. They support flexible frame sampling and facilitate evaluation under diverse temporal intervals.
  • Xiph [195, 60]: A curated set of 4K video sequences designed for assessing interpolation fidelity in subtle motion settings.
  • KITTI [190]: Captured in autonomous driving scenarios, KITTI poses unique challenges with sparse GT and large ego-motion.
@@
  • Sintel [192]: A synthetic dataset rendered from the Sintel film, offering photorealistic motion and structured flow annotations.
  • DAVIS [191]: Originally for segmentation, DAVIS features complex object motion, occlusion, and deformation, offering rich dynamics for interpolation.
@@
  • Adobe240 [179]: Collected at 240fps, this dataset captures real-world motion blur and lighting changes, ideal for fine-grained temporal modeling.
@@
  • GOPRO [193]: Featuring high-frame-rate recordings with handheld cameras, GOPRO provides realistic non-linear motion and defocus blur.
@@
  • HD [33]: A subset of HR content from Xiph, with sharper motion content suited for realistic evaluation.
  • X4K1000FPS [79]: A premier benchmark for ultra-slow motion and long-range interpolation, thanks to its dense 1000fps and 4K capture settings.
  • WebVid-10M [194]: A large-scale web video corpus originally built for text-video tasks. Its size and diversity support generative VFI when properly filtered.
@@
  • LAVIB [187]: Designed for large-scale, diverse-domain evaluation with balanced splits and curated subsets for out-of-distribution testing.
@@
  • OpenVid [186]: A text-video dataset supporting multi-modal VFI and DM-based interpolation research via dense, aligned samples.
5.2 Data Augmentation
@@
Modern VFI models incorporate spatial and temporal data augmentation to improve generalization and prevent overfitting. A widely adopted strategy is patch-based cropping, where fixed-size patches (e.g., $128\times 128$ or $256\times 256$) are randomly extracted from HR inputs [29, 62, 81, 54]. This not only reduces memory and computational costs but also encourages localized motion learning while mitigating spatial overfitting to scene layout or object positioning. Furthermore, random cropping prevents the model from overfitting to spatial priors such as background layout or object location, thereby improving robustness across spatial contexts [29]. Additional spatial augmentations, such as horizontal/vertical flipping and random rotation, enhance appearance diversity and promote invariance to orientation and perspective changes. These augmentations enable the model to remain invariant to directional biases and better generalize to unseen spatial transformations.
@@
Temporal augmentation is equally critical in sequential modeling. Frame order reversal [81, 54] is commonly applied, wherein sequences like $(I_{0},I_{1},I_{2})$ are reversed to $(I_{2},I_{1},I_{0})$. In CTFI, this augmentation preserves the center-frame $I_{1}$ while exposing the model to symmetric motion trajectories [5, 50]. Similarly, in ATFI settings, reversing sequences ensures temporal consistency under bidirectional motion. For example as shown in Fig. 9 (b), consider an input triplet $(I_{0},I_{\frac{1}{3}},I_{1})$ used to supervise interpolation at $t{=}\frac{1}{3}$. By reversing the sequence to $(I_{1},I_{\frac{1}{3}},I_{0})$, the relative time becomes $(1{-}\frac{1}{3}){=}\frac{2}{3}$. This simple yet effective strategy enables the model to learn temporally symmetric representations, thereby improving generalization across motion directions and enhancing robustness in bidirectional synthesis.
@@
TABLE II: Summary of evaluation metrics for VFI.
[…]
Category Metric Interpolation Quality Reference Frame
[…]
5.3 Evaluation Metrics
To facilitate comprehensive assessment of VFI models, various metrics have been proposed to capture different aspects of visual quality and temporal coherence. Table II summarizes commonly used evaluation metrics categorized into image-level, perceptual, and video-level types.
5.3.1 Image-level Metrics

### v3

@@ IV-B Occlusion
Achieving high-quality (HQ) interpolation demands accurate motion estimation as well as a proper understanding of occlusions. Otherwise, severe artifacts are likely to appear in the predicted frames, particularly near motion boundaries. For two consecutive input frames, certain pixels in the intermediate frame may not correspond to any observable region in either input, creating ambiguity in determining the correct content for these occluded regions [233]. As shown in Fig. 10 (b), such occlusions can occur when previously hidden objects become visible or when objects move toward the camera, revealing regions that were not seen in either input. Naively blending warped inputs often results in severe artifacts, most notably ghosting artifacts [104], where an object is not only incorrectly projected from its previous location but also appears as a duplicate at its correct position due to the lack of sufficient visual cues. This is especially problematic in disoccluded regions, areas newly revealed in the intermediate frame but absent in both inputs, such as when an object emerges from behind another or moves directly toward the viewpoint. In these cases, the absence of visual evidence introduces ambiguity, making it unclear what content should be synthesized. To resolve this, modern VFI methods incorporate explicit occlusion reasoning to guide the synthesis process.
@@
Overall, occlusion-aware VFI remains a critical challenge, particularly in dynamic scenes with depth discontinuities or disoccluded motion. As such, SOTA models increasingly combine multiple strategies, such as masking, depth priors, feature similarity, or forward-backward consistency [155, 104] to recover plausible content in ambiguous regions and maintain temporal coherence in the output. To complement the quantitative statistics, Fig. 11 (top row) presents a failure-case comparison on Vimeo90K [252], which contains significant occlusions [204]. Here the flow-based GIMM-VFI produces noticeable ghosting artifacts around disocclusion boundaries (purple arrows), while the kernel-based FLAVR yields smoother although slightly blurred transitions inside the occluded region (yellow arrows). The DM-based FCVG is able to hallucinate a plausible object shape behind the occluder, improving perceptual completeness but deviating mildly from the GT geometry (green arrows). Together with the Adobe240fps example, which lies between extreme occlusion and large-motion regimes, these qualitative behaviors highlight that kernel-based sampling can be relatively more stable in heavily occluded or texture-poor regions, DM-based models favor perceptual plausibility even when exact geometry is ambiguous, and flow-based models remain sensitive to flow estimation errors near occlusion boundaries.
IV-C Lighting Variation
@@
TABLE I: Summary and comparison of popular datasets for VFI. The dataset types
@@
represent Triplet datasets, and
@@
represent Multi-frame datasets.
@@
IV-D Non-linear Motion
Many early VFI methods [148, 107, 167, 7, 6, 145, 251, 262, 252, 177, 168, 109] assume linear or uniform motion between input frames. Under this assumption, objects move along straight trajectories at constant velocity, allowing motion estimation based on simple temporal interpolation. Flow-based [148, 107, 7, 252], kernel-based [7], and even phase-based models [158] often rely on this assumption implicitly. However, in real-world scenarios, motion is frequently non-linear due to acceleration, deceleration, directional changes, or complex local dynamics. As shown in Fig. 10 (d), a beach ball is thrown along a curved trajectory over a wavy sea surface; both the curved object path and the rapidly evolving wave patterns violate the linear-motion prior and lead to significant estimation errors when only two frames are observed. In practice, such non-linear behavior can be broadly observed in three forms: (i) curved or acceleration-driven object trajectories, (ii) fine-grained dynamic textures such as water or foliage, and (iii) velocity ambiguity, where multiple plausible velocity profiles or motion paths produce the same observed displacement between two frames.
To address the limitations of linear motion assumptions, researchers have proposed higher-order motion modeling that extends beyond simple first-order trajectories. Since most existing methods operate on only two input frames, they are inherently under-constrained and often forced to assume simple motion. To overcome this, several methods [251, 34, 195, 144] incorporate multiple input frames (typically four) to capture richer temporal variations and better approximate non-linear motion. QVI [251] introduces a quadratic motion model that fits second-order trajectories over four consecutive input frames. Specifically, it takes $(I_{-1},I_{0},I_{1},I_{2})$ as inputs and predicts an intermediate frame $I_{t}$ for arbitrary $t\in(0,1)$. By modeling both velocity and acceleration from surrounding frames, QVI enables the network to better handle curved or time-varying motion paths. This parametric formulation allows the model to explicitly account for motion curvature. EQVI [144] further refines this idea by combining offset-based warping with temporal embeddings, improving precision and robustness. More recently, IQ-VFI [88] introduces an implicit motion representation using a coordinate-based MLP that adapts to arbitrary motion patterns without requiring predefined trajectory assumptions. These works collectively emphasize the importance of modeling non-linear motion directly, especially in multi-frame settings. Nevertheless, explicitly parameterized motion models such as quadratic trajectories still cannot fully capture the complexity and irregularity of real-world motions, particularly under strong occlusions or highly non-rigid dynamics.
Non-linear motion is also prominent in dynamic textures, including water, fire, smoke, or foliage [46]. These regions contain high spatial-frequency content with temporally stochastic and spatially irregular motion, where local displacements are multi-directional and only weakly correlated over time. Conventional flow-based [107, 148] and kernel-based models [170] that assume smooth, one-to-one motion fields often produce temporal incoherence and “boiling” artifacts on such content. ST-MFNet [46] mitigates this by combining multi-flow warping with a 3D CNN-based texture enhancement branch, while TAFI [41] conditions on texture classes (static, dynamic continuous / discrete) to specialize interpolation behavior.
Finally, velocity ambiguity [281] has recently been recognized as another key challenge. When only two frames are available, multiple plausible motion trajectories can explain the observed displacement, making the underlying motion inherently under-constrained. This ambiguity becomes especially pronounced in scenes involving curved motion or directional switches, such as bouncing balls or rotating limbs. To tackle this, Zhong et al. [281] learns a velocity embedding that jointly reasons about motion direction and temporal consistency, while BiM-VFI [194] introduces Bidirectional Motion Fields (BiM) encoding angular and magnitude differences around the intermediate time to better model curved and asymmetric trajectories.
These recent advances collectively signal a shift from rigid linear motion priors toward flexible, context-aware motion modeling in VFI. By extending temporal supervision and refining motion representations, either through higher-order parametric formulations, implicit coordinate-based embeddings, or bidirectional velocity fields, modern VFI methods achieve significantly improved performance in complex non-linear motion scenarios. At the same time, such regimes often expose discrepancies between distortion-based metrics (e.g., PSNR) and perceptual or temporal-consistency measures, suggesting that fully understanding and benchmarking non-linear motion remains an open and practically important challenge for VFI.
V Datasets and Evaluation
V-A Datasets
To facilitate training and evaluation across varying temporal resolutions and motion complexities, numerous VFI datasets have been developed. Tab. I provides a high-level summary of commonly used datasets categorized into triplet and multi-frame types. We describe each dataset in detail below.
V-A1 Triplet Datasets
Early deep learning-based VFI approaches primarily rely on triplet datasets, where two input frames are used to predict the temporally centered GT frame. This configuration aligns with CTFI settings (Sec. III-A). Some datasets are further extended to seven-frame sequences [252] for evaluating frame-rate upsampling.
  • Middlebury [5]: Originally designed for optical flow, it contains short video clips with moderate complexity.
  • UCF101 [209, 148]: A human action dataset from which a small subset of triplets is used for VFI.
  • Vimeo90K [252]: A widely adopted benchmark with diverse scenes and consistent format. It offers clean supervision and balanced motion complexity.
  • SNU-FILM [36]: Constructed from high-speed footage and categorized by motion difficulty, it enables evaluation across varying levels of motion, occlusion, and blur.
  • ATD-12K [208]: A large-scale animation dataset with rich stylistic diversity.
V-A2 Multi-frame Datasets
@@
Multi-frame datasets enable dense temporal supervision and are commonly used in both CTFI and ATFI (Sec. III-B) settings. They support flexible frame sampling and facilitate evaluation under diverse temporal intervals.
  • Xiph [161, 168]: A curated 4K video suite with fine, predominantly small motions, widely used to assess interpolation fidelity under high-resolution, subtle-motion conditions.
  • KITTI [63]: Driving scenes captured from a moving vehicle, characterized by sparse ground truth, large ego-motion, and strong perspective changes, forming a challenging benchmark for motion estimation and interpolation.
@@
  • Sintel [18]: A synthetic dataset rendered from the Sintel film, providing photorealistic appearance, complex dynamics, and dense optical flow annotations for controlled evaluation.
  • DAVIS [184]: Originally proposed for video object segmentation, DAVIS features complex object motion, occlusions, and non-rigid deformations, making it suitable for analyzing interpolation behavior in realistic dynamic scenes.
@@
  • Adobe240 [211]: Real-world videos captured at 240 fps, containing motion blur and illumination changes, commonly used to evaluate fine-grained temporal modeling and slow-motion synthesis.
@@
  • GOPRO [163]: High-frame-rate recordings from handheld cameras with non-linear motion and defocus blur, providing a realistic testbed for VFI under camera shake and depth-varying blur.
@@
  • HD [7]: A high-resolution subset derived from Xiph, including sharper content and more pronounced motion, suitable for realistic HR interpolation evaluation.
  • X4K1000FPS [204]: A 4K, 1000 fps dataset designed for ultra–slow motion and long-range interpolation, offering dense temporal sampling for high-fidelity frame synthesis studies.
  • WebVid-10M [4]: A large-scale web video corpus originally curated for text–video learning; after appropriate filtering, it provides diverse, in-the-wild content for generative and data-driven VFI.
@@
  • LAVIB [210]: A large-scale, diverse-domain benchmark with balanced splits and curated subsets, enabling systematic evaluation of in-distribution and out-of-distribution VFI performance.
@@
  • OpenVid [165]: A text–video dataset with densely aligned samples that supports multi-modal VFI and diffusion-based interpolation research under rich semantic conditioning.
TABLE II: Summary of evaluation metrics for VFI.
Arrows ( / ) indicate whether higher or lower values correspond to better interpolation quality. A checkmark ( ) indicates that the metric requires GT frames. Colored rows denote perceptual metrics.
Category Metric Interpolation Quality Reference Frame
Image-level Metrics PSNR
SSIM [234]
IE [5]
NIQE [160]
FID [80]
LPIPS [273]
FloLPIPS [45]
STLPIPS [64]
DISTS [51]
Video-level Metrics VSFA [129]
tOF [37]
FVD [223]
FVMD [142]
VBench [95]
V-B Data Augmentation
@@
Modern VFI models incorporate spatial and temporal data augmentation to improve generalization and prevent overfitting. A widely adopted strategy is patch-based cropping, where fixed-size patches (e.g., $128\times 128$) are randomly extracted from HR inputs [170, 34, 120, 133]. This not only reduces memory and computational costs but also encourages localized motion learning while mitigating spatial overfitting to scene layout or object positioning. Furthermore, random cropping prevents the model from overfitting to spatial priors such as background layout or object location, thereby improving robustness across spatial contexts [170]. Additional spatial augmentations, such as horizontal or vertical flipping and random rotation, enhance appearance diversity and promote invariance to orientation and perspective changes. These augmentations enable the model to remain invariant to directional biases and better generalize to unseen spatial transformations.
@@
Temporal augmentation is equally critical in sequential modeling. Frame order reversal [120, 133] is commonly applied, wherein sequences like $(I_{0},I_{1},I_{2})$ are reversed to $(I_{2},I_{1},I_{0})$. In CTFI, this augmentation preserves the center-frame $I_{1}$ while exposing the model to symmetric motion trajectories [107, 36]. Similarly in ATFI settings, reversing sequences ensures temporal consistency under bidirectional motion. For example as shown in Fig. 9 (b), consider an input triplet $(I_{0},I_{\frac{1}{3}},I_{1})$ used to supervise interpolation at $t{=}\frac{1}{3}$. By reversing the sequence to $(I_{1},I_{\frac{1}{3}},I_{0})$, the relative time becomes $(1{-}\frac{1}{3}){=}\frac{2}{3}$. This simple yet effective strategy enables the model to learn temporally symmetric representations, thereby improving generalization across motion directions and enhancing robustness in bidirectional synthesis.
@@
TABLE III: Quantitative comparison on standard VFI benchmarks. We summarize the performance of representative methods on Vimeo-90K [252], Xiph-2K/4K [168], and SNU-FILM (Extreme) [36]. The results are reported in the order of PSNR $\uparrow$ / LPIPS $\downarrow$.
Best viewed in zoom.
Methods Vimeo-90K Xiph-4K Xiph-2K SNU-FILM (Extreme)
Kernel-based (§II-B1)
SepConv [ICCV’17] [170] 33.790 / 0.027 32.060 / 0.169 34.770 / 0.067 24.653 / 0.183
Flow-based (§II-B2)
DAIN [CVPR’19] [6] 34.700 / 0.022 33.490 / 0.170 35.950 / 0.084 24.819 / 0.142
SoftSplat [CVPR’20] [168] 36.100 / 0.021 33.600 / 0.234 36.620 / 0.107 25.436 / 0.119
ABME [ICCV’21] [178] 36.180 / 0.021 33.730 / 0.236 36.530 / 0.107 25.420 / 0.182
XVFI [ICCV’21] [204] 35.070 / 0.023 32.450 / 0.184 35.170 / 0.084 24.677 / 0.139
RIFE [ECCV’22] [93] 34.160 / 0.020 33.760 / 0.207 36.190 / 0.092 24.840 / 0.139
IFRNet [CVPR’22] [120] 36.200 / 0.019 33.970 / 0.136 36.570 / 0.068 25.270 / 0.116
FILM [ECCV’22] [188] 35.710 / 0.013 33.830 / 0.184 36.530 / 0.091 25.170 / 0.106
AMT [CVPR’23] [133] 35.790 / 0.021 34.653 / 0.199 36.415 / 0.089 25.430 / 0.112
UPR-Net [CVPR’23] [109] 36.420 / 0.020 33.647 / 0.230 36.749 / 0.103 25.630 / 0.112
SGM-VFI [CVPR’24] [139] 35.810 / 0.023 33.260 / 0.221 36.060 / 0.101 25.380 / 0.118
Transformer-based (§II-B6)
CAIN [AAAI’20] [36] 34.650 / 0.020 32.560 / 0.223 35.210 / 0.103 25.060 / 0.203
VFIFormer [CVPR’22] [150] 36.380 / 0.021 33.370 / 0.227 36.550 / 0.107 25.430 / 0.119
EMA-VFI [CVPR’23] [270] 36.340 / 0.026 33.260 / 0.219 36.540 / 0.097 25.690 / 0.114
Mamba-based (§II-B7)
VFIMamba [NeurIPS’24] [267] 36.090 / 0.021 34.260 / 0.218 36.710 / 0.101 25.590 / 0.059
LC-Mamba [CVPR’25] [105] 36.190 / 0.022 34.260 / 0.214 36.670 / 0.100 25.330 / 0.060
V-C Evaluation Metrics
To facilitate comprehensive assessment of VFI models, various metrics have been proposed to capture different aspects of visual quality and temporal coherence. Tab. II summarizes commonly used evaluation metrics categorized into image-level, perceptual, and video-level types.
V-C1 Image-level Metrics